Teal Company has been operating for several years, and on December 31, 2020, presented the following balance sheet.

TEAL COMPANY
BALANCE SHEET
DECEMBER 31, 2020

Cash

$39,100

Accounts payable

$72,400

Receivables

79,400

Mortgage payable

143,000

Inventory

87,500

Common stock ($1 par)

147,300

Plant assets (net)

225,400

Retained earnings

68,700
$431,400 $431,400


The net income for 2020 was $25,500. Assume that total assets are the same in 2019 and 2020.

(a) Compute each of the following ratios.

1) Current ratio

2) Acid-test ratio

3) Debt to assets ratio

4) Return on assets
Answer #1

Current ratio = Current Assets / Current Liabilities

Current Assets = Cash + Receivables + Inventory

=39100 + 79400 +87500

= 206000

Current Liabilities = Accounts Payable = 72400

Current ratio = 206000 / 72400 = 2.845

Acid-test ratio = Current Assets - Inventories / Current Liabilities

206000 - 87500 = 118500

Current Liabilities = Accounts Payable = 72400

Acid-test ratio = 118500 / 72400

1.636

Debt to assets ratio = Short term debt + long term debt / Total assets

Short term debt + long term debt = 72400 + 143000 = 215400

Total assets = 431400

Debt to assets ratio = 215400 / 431400 = 0.499

Return on assets = Net Income / Average Total assets

Net Income = 25500

Average Total assets = opening assets + closing assets / 2

=431400 + 431400 /2

= 431400

Return on assets = 25500 / 431400 x 100

5.91%

Mortgage Payable is considered to be long term debts hence not taken in calculation of Current Ratio and Acid test ratio